# On an approach to canonicalizing elliptic Feynman integrals

### v2

@@ B.2.2 The construction with constant period matrix
$\phi_{i-3}\to\big{(}\,\mathcal{T}_{1}\cdot\vec{\phi}\,\big{)}_{i-3}=\phi_{i-3}-E(\sin^{-1}a_{i},k^{2})\phi_{1}+F(\sin^{-1}a_{i},k^{2})\phi_{n-1}\,,$ (129)
with $i=5,\cdots,n$. The transformation is essentially the same as Eq. 105, and the only difference arises in different choice of pre-canonical basis integral for the second kind, which is immaterial. Next we only need to focus on the block for $\phi_{1}$ and $\phi_{n-1}$, whose period matrix is
@@ B.2.3 From Jacobi family back to the original family
$\phi_{1}$ $=\frac{c_{41}}{2}\frac{{\rm d}z}{\sqrt{P_{4}(z)}}\,,$ (143a)
$\phi_{i-3}$ $=\frac{1}{z-c_{i}}\frac{\sqrt{P_{4}(c_{i})}}{\sqrt{P_{4}(z)}}{\rm d}z\,,$ (143b)
$\phi_{n-2}$ $=\left(z-\frac{c_{2}+c_{3}}{2}\right)\frac{{\rm d}z}{\sqrt{P_{4}(z)}}\,,$ (143c)
$\phi_{n-1}$ $=\frac{2}{c_{14}}(z-c_{1})(z-c_{4})\frac{{\rm d}z}{\sqrt{P_{4}(z)}}\,,$ (143d)
@@
We want to emphasize that the example mentioned above represents a very special case, where the point at infinity in the Jacobi normal form corresponds directly to the point at infinity in the original form. However, in more general cases, the point at infinity in the normal form may differ from the point at infinity in the original form. This distinction complicates the transformation process, but it can be circumvented by employing an alternative representation for the elliptic integral of the second kind, e.g.
$\phi_{n-1}=\frac{1-k^{2}}{k}\frac{1}{t+\frac{1}{k}}\frac{{\rm d}t}{\sqrt{P_{J}(t)}}\,,$ (144)
and other differentials with a double pole at $t=-1,+1,+1/k$ are also possible. While for the ${\rm d}\log$-form, we replace it with the elliptic integral of the third kind
$\phi_{n-2}=\frac{1}{t-a_{\infty}}\frac{\sqrt{P_{J}(a_{\infty})}}{\sqrt{P_{J}(t)}}{\rm d}t\,,$ (145)
@@
where the $k$ specifying the Jacobi curve here is the modulus $k^{2}=\lambda$. We can see the twist of the Jacobi family is in a more restricted form, where the marked points ±$\pm a_{i}$ appear in pairs. Moreover, the number of the basis integrals is twice as many as the original. However, we can always construct the additional pre-canonical basis integrals as ${\rm d}\log$-forms, which disentangles with other ones in the $\varepsilon^{0}$-part of the differential equations. However, the ${\rm d}\log$-forms cannot be transformed back to the original family, while the other ones belong to the original family so they can indeed serve as pre-canonical basis integrals in the original family.
@@
B.3 Application to the non-planar double box family
Although the simple pre-canonical basis integrals in Eqs. (143) ask a special original family, it turns out our first non-planar double box family in Sec. 3.2 satisfies the condition. It is important to note that the underlying elliptic curve possesses a unique property: the branch points exhibit an additional symmetry. We can refer back to the branch points Eqs. (80). We could apply the results above directly. In this case, the corresponding Möbius transformation is
$z\mapsto t=S(z)=\frac{2z-1}{\sqrt{1-16y_{s}}}\,,$ (148)
and its reverse is
$z=S^{-1}(t)=\frac{1}{2}\left(\sqrt{1-16y_{s}}t+1\right)\,.$ (149)
Under the transformation, the branch points $c_{i}$’s are mapped to $a_{i}$’s , and we can easily know the parameter $k$ for the elliptic curve in the Jacobi normal form with Eq. (150). It should be
$k=\sqrt{1-16y_{s}}\,.$ (150)
We follow the same steps in Sec. 3.2. The explicit form of the pre-canonical basis in the original integral family Eqs. (143) should be
$\phi_{1}^{\rm npdb}$ $=\frac{1}{2}\frac{{\rm d}z}{\sqrt{z(z-1)[z(z-1)+4y_{s}]}}\,,$ (151a)
$\phi_{2}^{\rm npdb}$ $=\frac{1}{z-y_{t}}\frac{\sqrt{(y_{t}-1)y_{t}[4y_{s}+(y_{t}-1)y_{t}]}}{\sqrt{z(z-1)[z(z-1)+4y_{s}]}}{\rm d}z\,,$ (151b)
$\phi_{3}^{\rm npdb}$ $=\frac{1}{2}(2z-1)\frac{{\rm d}z}{\sqrt{z(z-1)[z(z-1)+4y_{s}]}}\,,$ (151c)
$\phi_{4}^{\rm npdb}$ $=-2z(z-1)\frac{{\rm d}z}{\sqrt{z(z-1)[z(z-1)+4y_{s}]}}\,.$ (151d)
We could see this basis enjoys a simpler form than the one in Eqs. (99), and here algebraic factors only appear in $\phi_{2}$. We also note that the pre-canonical basis integrals in Becchetti:2025rrz are very similar to the ones above.
For the explicit form, $k$ is given by Eq. (150) and $a_{5}$ is given by
$a_{5}=\frac{2y_{t}-1}{\sqrt{1-16y_{s}}}\,.$ (152)
@@ B.3 Application to the non-planar double box family
In order to reach a canonical basis for the whole family, we only need to apply essentially the same strategy in Sec. 3.2. One can check the two constructions are equivalent. However, the involved elliptic integrals for Eqs. (151) and the corresponding transformation to the canonical basis have simpler arguments than the ones for the basis in Sec. 3.2. The elliptic integrals with distinct arguments can be related by nontrivial identities like the ones in App. D.
@@ Appendix C The pre-canonical basis in the explicit form
In this appendix, we will provide the explicit form of the pre-canonical basis for which the differential equation is $\varepsilon$-factorized, except in the top-sector. To achieve a fully $\varepsilon$-factorized differential equation, one can apply the transformation from (39) within the top-sector.
@@
$\left.-4y_{s}\left(2\left(4\varepsilon\left(\sqrt{1-16y_{s}}-1\right)+\sqrt{1-16y_{s}}\right)I_{11111110(-1)}\right.\right.$
$\left.\left.+\left(\sqrt{1-16y_{s}}-1\right)\left((16y_{s}-1)I_{111210100}+(1-8y_{s})I_{11121110(-1)}-8y_{s}I_{21111110(-1)}\right.\right.\right.$
@@
$\left.+2\left((1-\varepsilon)y_{s}I_{121101000}-\varepsilon\left(4y_{s}+y_{t}-1\right)I_{201002000}\right.\right.$
$\left.\left.+\varepsilon\left((2\varepsilon-1)y_{s}+(\varepsilon-1)y_{t}\right)I_{110201000}\right.\right.$
@@ Appendix D Special relations of elliptic integrals
With Eqs. (136) and (8), we obtain a quadratic transformation which transform a standard elliptic integral with parameter $k^{2}$ to the one with parameter $\lambda=4k/(1+k)^{2}$. The transformation is given by

### v3

@@ B.2.2 The construction of an $\varepsilon$-factorized basis with constant period matrix
$\phi_{i-3}^{\rm pre}\to\big(\,\mathcal{T}_{1}\,\vec{\phi}^{\rm pre}\,\big)_{i-3}=\phi_{i-3}^{\rm pre}-E(\sin^{-1}a_{i},k^{2})\phi_{1}^{\rm pre}+F(\sin^{-1}a_{i},k^{2})\phi_{n-1}^{\rm pre}\,$, (140)
with $i=5,\cdots,n$. The transformation is essentially the same as Eq. 116, and the only difference arises in different choice of pre-canonical basis integral for the second kind, which is immaterial. Next we only need to focus on the block for $\phi_{1}^{\rm pre}$ and $\phi_{n-1}^{\rm pre}$, whose period matrix is
@@ B.2.3 From Jacobi family back to the original family
$\phi_{1}^{\rm pre}$ $=\frac{c_{41}}{2}\frac{{\rm d}z}{\sqrt{P_{4}(z)}}\,$, (154a)
$\phi_{i-3}^{\rm pre}$ $=\frac{1}{z-c_{i}}\frac{\sqrt{P_{4}(c_{i})}}{\sqrt{P_{4}(z)}}{\rm d}z\,$, (154b)
$\phi_{n-2}^{\rm pre}$ $=\left(z-\frac{c_{2}+c_{3}}{2}\right)\frac{{\rm d}z}{\sqrt{P_{4}(z)}}\,$, (154c)
$\phi_{n-1}^{\rm pre}$ $=\frac{2}{c_{14}}(z-c_{1})(z-c_{4})\frac{{\rm d}z}{\sqrt{P_{4}(z)}}\,$, (154d)
@@
We want to emphasize that the example mentioned above represents a very special case, where the point at infinity in the Jacobi normal form corresponds directly to the point at infinity in the original form. However, in more general cases, the point at infinity in the normal form may differ from the point at infinity in the original form. This distinction complicates the transformation process, but it can be circumvented by employing an alternative representation for the elliptic integral of the second kind, e.g.,
$\phi_{n-1}^{\rm pre}=\frac{1-k^{2}}{k}\frac{1}{t+\frac{1}{k}}\frac{{\rm d}t}{\sqrt{P_{J}(t)}}\,$, (155)
and other differentials with a double pole at t=1,+1,+1/kt=-1,+1,+1/k are also possible. While for the ${\rm d}\log$-form, we replace it with the elliptic integral of the third kind
$\phi_{n-2}^{\rm pre}=\frac{1}{t-a_{\infty}}\frac{\sqrt{P_{J}(a_{\infty})}}{\sqrt{P_{J}(t)}}{\rm d}t\,$, (156)
@@
where the $k$ specifying the Jacobi curve here is the modulus $k^{2}=\lambda$. We can see the twist of the Jacobi family is in a more restricted form, where the marked points $\pm a_{i}$ appear in pairs. Moreover, the number of the basis integrals is twice as many as the original. However, we can always construct the additional pre-canonical basis integrals as ${\rm d}\log$-forms, which disentangles with other ones in the $\varepsilon^{0}$-part of the differential equations. However, the ${\rm d}\log$-forms cannot be transformed back to the original family, while the other ones belong to the original family so they can indeed serve as pre-canonical basis integrals in the original family.
@@
B.3 Application to non-planar double box family
Although the simple pre-canonical basis integrals in Eqs. (154) ask a special original family, it turns out our first non-planar double box family in Sec. 3.4 satisfies the condition. It is important to note that the underlying elliptic curve possesses a unique property: the branch points exhibit an additional symmetry. This symmetry makes the construction in Jacobi form simpler to use.
For the branch points of the single mass case in Sec. 3.4, we can refer back to Eqs. (85). With the Möbius transformation in Eq. (149), the variables $k$ and $a_{5}$ for the elliptic curve in the Jacobi normal form are
$k=\sqrt{1-16y_{s}}\,,\quad a_{5}=\frac{2y_{t}-1}{\sqrt{1-16y_{s}}}\,.$ (159)
Plugging Eqs. (159) into Eqs. (125), one can check this pre-canonical basis enjoys an arguably simpler form than the one obtained in Sec. 3.4, and here algebraic factors only appear in $\phi_{2}^{\rm pre}$. We also note that the pre-canonical basis integrals in Becchetti:2025rrz are very similar to the ones here.
@@ B.3 Application to non-planar double box family
In order to reach a canonical basis for the whole family, we only need to apply essentially the same strategy in Sec. 3.4. One can check the two constructions are equivalent. However, the involved pre-canonical basis and ${\cal T}_{\rm can}$ have simpler arguments than the ones for the basis in Sec. 3.4. The elliptic integrals with distinct arguments can be related by non-trivial identities like the ones in App. D.
@@ Appendix C The pre-canonical basis in the explicit form
In this appendix, we will provide the explicit form of the pre-canonical basis of Sec. 3.4 for which the differential equation is $\varepsilon$-factorized, except in the top-sector. To achieve a fully $\varepsilon$-factorized differential equation, one can apply the transformation from (25) within the top-sector.
@@
[…]
@@
[…]
@@ Appendix D Special relations of elliptic integrals
With Eqs. (147) and (17), we obtain a quadratic transformation which transforms a standard elliptic integral with parameter $k^{2}$ to the one with parameter $\lambda=4k/(1+k)^{2}$. The transformation is given by